The specific weight of glycerin is 78.6 lb/ft3 . Compute its density and specific gravity. What is its specific weight in kN/m3 ?


1
Expert's answer
2022-03-02T17:00:03-0500

"\rho=1260~\frac{kg}{m^3},"

"S_g=\frac{\rho}{\rho_0}=1.26,"

"\gamma=\rho g=12.36~\frac{kN}{m^3}."